# A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance

### v4

@@ Abstract
This paper presents a volumetric privacy framework for dynamical systems subject to bounded disturbances, developed without requiring prior knowledge of their probability distributions. We consider systems with both public and private states, where a set containing the public state is shared as the observation. An adversary is assumed to execute an inference attack by exploiting the observed public state set to estimate an uncertainty set for the private state. The volume of this inferred set quantifies the adversary’s estimation uncertainty and serves as the proposed volumetric privacy metric. Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that it is bounded by the information gain from the observation set. Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the volumetric privacy level. The effectiveness of the proposed approach is validated through a production–inventory case study. Results show that the optimal privacy filter significantly improves robustness against inference attacks and outperforms two baseline mechanisms based on additive noise and quantization.
@@ 1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational information can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring privacy protection for dynamic systems.
Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information, e.g., inventory–production systems and connected vehicles. However, the notion of privacy in such systems remains insufficiently explored. Within the context of set-membership estimation, the states of these systems can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance. We develop privacy-preserving strategies aimed at maximizing an adversary’s uncertainty, i.e., the volume of uncertainty set, in inferring private states. The proposed approaches are applicable to both deterministic and stochastic systems, without requiring prior knowledge of the underlying probability distributions.
@@ 1.2 Related Work
Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data. While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs. However, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation. Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec. 5.
Finally, other deterministic privacy-preserving approaches often exploit observability reduction or state decomposition to protect private information in multi-agent systems. For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states. In [3], the trace of the observability Gramian was employed to quantify information leakage through an intruder node, and an online optimization strategy was proposed to adapt communications in order to minimize such leakage. Privacy-preserving consensus designs were also proposed in [25, 26], including augmented states and novel consensus algorithms with simultaneous accuracy, resilience, and privacy guarantees. State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
@@ 1.3 Contributions
This paper studies privacy for dynamic systems with bounded disturbance as shown in Fig. 1, where the system state is split into public $X_{k}$ and private $Y_{k}$, and an adversary uses the public observation set $\mathcal{M}_{k|k}^{x}$ to infer $Y_{k}$ via an uncertainty set $\mathcal{Y}_{k|k}$.
@@ 1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model, the inference attack and defines the volumetric privacy and utility measure. Section 3 provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
@@ 1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. For the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z\!+\!R|Z\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$ with the absolute value $\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The 1-norm of the matrix $A$ is defined as $\left\|A\right\|_{1}\!\frac{\Delta}{=}\!\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$. The operator $diag(v)$ denotes a diagonal matrix constructed from the vector $v$.
2 System Model and Inference Attack
@@ 2.1 System Model
We consider the following stable system model $\mathbf{G_{1}}$,
$\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right$., (3)
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the unknown disturbance with bounded sets $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$. Note that the underlying probability distributions of $W_{k}^{x}$ and $W_{k}^{y}$ are unknown. Besides, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. We assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
@@ 2.2 Motivating Examples
In this subsection, we consider two motivating examples to illustrate the necessity of protecting privacy of systems $\mathbf{G_{1}}$.
@@
2.3 Inference Attack
Due to the presence of unknown disturbance terms belonging to bounded sets, multiple private states may correspond to the same public state. Consequently, for a given public state set, an adversary cannot determine the exact private state; instead, it can only identify a corresponding uncertainty set. we next define how the adversary infers the private state using set-theoretic operations.
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$ and other elements to obfuscate the adversary’s estimation of private states. Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
@@ 2.3 Inference Attack
2.4 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. The attacker has more inference uncertainty about the private state if $\mathcal{Y}_{k|k}$ contains more elements. However, since the state space is continuous, the number of elements in such sets is uncountable. To address this, we propose to use the volume of the uncertainty set as a quantitative measure of privacy. Specifically, we define
where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure (i.e., the geometric volume) of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n}$. We then define the privacy measure at time $k$ as
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. To protect the system from inference attack, we can increase the privacy level of the private state via maximizing $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
On the other hand, it causes more utility distortion of the public state if the observation set $\mathcal{M}^{x}_{k|k}$ contains more elements, since it is more difficult for the receiver to recover the actual public state. To address this, we would reduce distortion via maximizing the following defined utility of public state set $\mathcal{M}^{x}_{k|k}$,
$U_{k}\left(\mathcal{M}^{x}_{k|k}\right)=1/\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right).$ (15)
Notably, set operations in the proposed inference attack and the volume computation are generally computational costly due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Also, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage. To address these challenges, we study computation approaches for inference attack in Sec. 3, and present an optimal privacy filter design to reduce privacy leakage in Sec. 4.
3 Inference Attack Approximation
@@ 3.1 Inference Attack Approximation via CCGs
$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$ (16)
@@ 3.2 Inference Attack Approximation via Interval Analysis
We next consider an interval-based approximation approach for computing the inference sets. This approach is a special case of CCG-based inference but significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic. Moreover, interval-based inference serves as the foundation for designing the optimal privacy filter, as discussed in Sec. 4.
Definition 5.
The interval $\mathcal{X}=\{X\mid\underline{X}\leqslant X\leqslant\overline{X}\}$ is defined as $\begin{bmatrix}\underline{X}\\[3.00003pt] \overline{X}\end{bmatrix}$ with the lower and upper bounds $\underline{X}$ and $\overline{X}$. An interval $\mathcal{X}$ can also be expressed as a special case of the CCG representation $\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathcal{R}^{n_{x}},\;\|\xi\|_{\infty}\leqslant 1\right\}$, with the center point $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and the radius $p^{x}=\frac{\overline{X}-\underline{X}}{2}$. Also, the volume of the interval $\mathcal{X}$ is computed as $\mathrm{Vol}\left(\mathcal{X}\right)=\sum_{i=1}^{n}{\left(\overline{X}\left(i\right)-\underline{X}\left(i\right)\right)}$ where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension of the interval $\mathcal{X}$, respectively.
Given a block matrix $A=[A_{1},\,A_{2}]$, the multiplication of $A$ with an interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. If $\mathcal{X}$ and $\mathcal{Y}$ are intervals, their Minkowski sum is $\mathcal{X}\oplus\mathcal{Y}=\begin{bmatrix}\underline{X}+\underline{Y}\\[3.00003pt] \overline{X}+\overline{Y}\end{bmatrix}$, and their difference, used only for volume evaluation, is $\mathcal{X}\setminus\mathcal{Y}=\begin{bmatrix}\underline{X}-\underline{Y}\\[3.00003pt] \overline{X}-\overline{Y}\end{bmatrix}$. The intersection of $\mathcal{X}$ and $\mathcal{Y}$ is $\mathcal{X}\cap\mathcal{Y}=\begin{bmatrix}\max\{\underline{X},\underline{Y}\}\\[3.00003pt] \min\{\overline{X},\overline{Y}\}\end{bmatrix}$.

### v5

@@ Abstract
In this paper, we first present a volumetric privacy measure for dynamical systems with bounded disturbances, wherein the states of the system contain private information and an adversary with access to sensor measurements attempts to infer the set of potential values of the private information. Under the proposed privacy measure, the volume of the uncertainty set of the adversary given the sensor measurements is considered as the privacy level of the system. We next characteristic the time evolution of the proposed privacy measure and study its properties for a particular system with both public and private states, where a set containing the public state is shared as the observation. Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that the uncertainty reduction of the adversary is bounded by the information gain from the observation set. Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the privacy level. The effectiveness of the proposed approach is validated through a production–inventory case study. Results show that the optimal privacy filter significantly improves robustness against inference attacks and outperforms two baseline mechanisms based on additive noise and quantization.
@@ 1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information[1]. Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information. Despite their importance, the notion of privacy in such systems remains insufficiently explored.
Within the context of set-membership estimation, the states of systems with bounded disturbance can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty. Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance. We develop privacy-preserving strategies aimed at maximizing an adversary’s uncertainty, i.e., the volume of uncertainty set, in inferring private states. The proposed approaches are applicable to both deterministic and stochastic systems, without requiring prior knowledge of the underlying probability distributions.
@@ 1.2 Related Work
Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data. While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation. Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec. 5.
Finally, other deterministic privacy-preserving approaches often exploit observability reduction or state decomposition to protect private information in multi-agent systems. For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states. The authors in [25, 26] presented privacy-preserving consensus algorithms that incorporate augmented states and novel consensus mechanisms, balancing accuracy, resilience, and privacy guarantees simultaneously. State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
@@ 1.3 Contributions
In this paper, we investigate volumetric privacy in dynamical systems subject to bounded disturbances, as illustrated in Fig. 1. The system state $S_{k}$ contains both utility information $X_{k}$ and private information $Y_{k}$, while an adversary exploits the observation set $\mathcal{M}_{k|k}^{x}$ to infer the private information $Y_{k}$ via the associated uncertainty set $\mathcal{Y}_{k|k}$.
@@ 1.4 Outline
The remainder of the paper is organized as follows. Section 2 introduces the system model and defines the volumetric privacy and utility measures. Section 3 formalizes the inference attack, presents computational approaches for evaluating the privacy level, and discusses the properties of the proposed measure. Section 4 develops an optimal privacy filter to mitigate privacy leakage while preserving a desired utility level. Numerical results are presented in Section 5, followed by concluding remarks in Section 6.
@@ 1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. For the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z\!+\!R|Z\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$ with the absolute value $\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The 1-norm of the matrix $A$ is defined as $\left\|A\right\|_{1}\!\frac{\Delta}{=}\!\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$. The operator $\mathrm{diag}(v)$ denotes a diagonal matrix constructed from the vector $v$.
2 System Model and Privacy Measure
@@ 2.1 System model
Consider the following stable dynamical system
$\mathbf{G}:\left\{\begin{array}[]{l}S_{k}=f\left(S_{k-1},W_{k}\right),\\[3.0pt] O_{k}=g\left(S_{k},V_{k}\right),\end{array}\right$. (3)
where $S_{k}$ denotes the system state, with initial condition $S_{0}$ belonging to a bounded set $\mathcal{S}_{0}\subset\mathcal{R}^{s}$, $O_{k}$ is the sensor measurement. The unknown disturbances $W_{k}$ and $V_{k}$ are assumed to be bounded within the sets $\mathcal{W}_{k}\subset\mathcal{R}^{s}$ and $\mathcal{V}_{k}\subset\mathcal{R}^{o}$, respectively. The system state $S_{k}$ contains both private and utility-related information, which are represented as continuous variables
$Y_{k}=h(S_{k}),\qquad X_{k}=u(S_{k}),$ (4)
where $Y_{k}$ represents the private information that must be kept confidential, and $X_{k}$ corresponds to the utility information intended to be disclosed. We assume that the adversary has full knowledge of system model $\mathbf{G}$ and will collect observations of the system to infer the private information.
A special case of this model is the following linear non-Gaussian system with invertible $A_{1}$ and $A_{2}$,
$\mathbf{G}_{1}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ O_{k}=X_{k}\\ \end{array}\right$., (8)
where the system state $S_{k}=[X_{k}^{\top},\,Y_{k}^{\top}]^{\top}$ consists of the public state $X_{k}\in\mathcal{R}^{n_{x}}$ and the private state $Y_{k}\in\mathcal{R}^{n_{x}}$, and the measurement $O_{k}$ corresponds to the public state $X_{k}$. The initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively.
@@ 2.2 Motivating Examples
We next consider two motivating examples to illustrate the necessity of protecting privacy of systems with bounded disturbance.
@@
2.3 Privacy and Utility Measures
Notably, in differential privacy for dynamical systems, the ranges of utility and private information are typically assumed to be time-invariant and unbounded, and privacy is characterized by probabilistic indistinguishability. However, for systems subject to bounded disturbances, the utility and private information can be represented by bounded uncertainty sets that are updated online using set-membership estimation. Accordingly, it is desirable to introduce a privacy measure tailored to such set-based descriptions.
Given the sensor measurements, an adversary can employ set-membership estimation techniques, e.g., [18, 19], to estimate both the utility and private information by constructing uncertainty sets $\mathcal{X}_{k|k}$ and $\mathcal{Y}_{k|k}$ that contain the true values $X_{k}=x_{k}$ and $Y_{k}=y_{k}$, respectively. We would further provide numerical approaches to illustrate how the adversary estimates $\mathcal{X}_{k|k}$ and $\mathcal{Y}_{k|k}$ in next section. This inference procedure, referred to as an inference attack, results in privacy leakage, as the adversary’s confidence in the private information increases when the uncertainty set $\mathcal{Y}_{k|k}$ shrinks. In the extreme case where $\mathcal{Y}_{k|k}=\{y_{k}\}$, the private value is fully revealed. Since $Y_{k}$ resides in a continuous space, the cardinality of $\mathcal{Y}_{k|k}$ is not meaningful, motivating a geometric approach to quantify uncertainty.
We therefore employ the volume of the uncertainty set as a quantitative measure of privacy:
where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n_{y}}$. The privacy measure at time $k$ is then defined as
A smaller volume corresponds to reduced privacy, while $\mathrm{Vol}(\mathcal{Y}_{k|k}\!)\!=\!0$ indicates complete exposure.
Similarly, the utility information can be represented by the uncertainty set $\mathcal{X}_{k|k}$. Larger $\mathcal{X}_{k|k}$ implies greater uncertainty in recovering the true public information $X_{k}=x_{k}$, resulting in higher distortion. Accordingly, we define the utility measure as
$U_{k}\!\left(\mathcal{X}_{k|k}\right):=\frac{1}{\mathrm{Vol}\!\left(\mathcal{X}_{k|k}\right)},$ (11)
so that higher $U_{k}$ corresponds to better utility. Throughout, we assume that all relevant uncertainty sets are measurable with finite, positive volume.
In the following, we analyze the proposed volumetric privacy measure in the context of the linear non-Gaussian system $\mathbf{G}_{1}$, and we develop practical approximations for evaluating the resulting privacy leakage. Based on this analysis, we design a privacy filter that mitigates leakage while satisfying utility requirements. A extension study of volumetric privacy for general nonlinear systems is left for future work.
3 Inference Attack for Linear Systems
3.1 Inference Attack via Set Operations
We assume that the adversary observes a set of public states, denoted by $\mathcal{M}_{k|k}^{x}$, which is generated by the privacy filter, as shown in Fig. 1. The observation set contains the true public state $X_{k}=x_{k}$ along with additional elements intended to obfuscate the adversary’s estimation of the private states. The adversary then performs the inference attack by identifying all possible values of the private state that are consistent with $\mathcal{M}_{k|k}^{x}$, thereby constructing its uncertainty set. In the following, we define the inference attack recursively.
@@ 3.2 Inference Attack Approximation via CCGs
$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$
@@ 3.3 Inference Attack Approximation via Interval Analysis
We next consider an interval-based approximation approach for computing the inference sets. This approach can be viewed as a special case of the CCG-based inference, but it significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic.
Definition 5 (Interval Representation).
An interval $\mathcal{X}=\{X\mid\underline{X}\leq X\leq\overline{X}\}$ is equivalently represented as $\mathcal{X}=\left[\begin{matrix}\underline{X}^{\top}&\overline{X}^{\top}\\ \end{matrix}\right]^{\top}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper bounds, respectively. Equivalently, an interval can be expressed as a special case of the CCG representation,
$\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathbb{R}^{n_{x}},\;\|\xi\|_{\infty}\leq 1\right\},$
with center $c^{x}=(\overline{X}+\underline{X})/2$ and radius $p^{x}=(\overline{X}-\underline{X})/2$. The volume of $\mathcal{X}$ is given by $\mathrm{Vol}(\mathcal{X})=\prod_{i=1}^{n_{x}}\big(\overline{X}(i)-\underline{X}(i)\big)=\prod_{i=1}^{n_{x}}2p^{x}(i)$, where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension, respectively.
To reduce computational complexity, a surrogate measure for the size of $\mathcal{X}$ can be defined as the total length across all dimensions $\overline{\mathrm{Vol}}(\mathcal{X})=\sum_{i=1}^{n_{x}}\big(\overline{X}(i)-\underline{X}(i)\big)=\sum_{i=1}^{n_{x}}2p^{x}(i)$, which is linear in the upper and lower bounds and therefore easier to compute. By the inequality of arithmetic and geometric means, this surrogate measure bounds the geometric volume, i.e., $\overline{\mathrm{Vol}}(\mathcal{X})/n\geq\sqrt[n]{\mathrm{Vol}(\mathcal{X})}$. Consequently, maintaining a large $\mathcal{X}$ implies a correspondingly large $\overline{\mathrm{Vol}}(\mathcal{X})$, making it a suitable surrogate metric.
[…]